OpenStudy (pphalke):

how many ounces of a 25% alcohol solution must be mixed with 11 ounces of a 30% alcohol solution to make a 26% alcohol solution?

OpenStudy (anonymous):

We have 11 oz of 30% alcohol We want to add some 25% alcohol to make 26% alcohol 0.25*x+0.3*11=0.26(11+x) 0.25*x+3.3=2.86+0.26x 0.25x-0.26x=2.86-3.3 -0.01x=-0.44 x=44 oz at 25% (11) is added to 11 oz at 30%(3.3) to make 26% of 55 oz(14.3) check 0.25*44+0.3*11=0.26(11+44) 11+3.3=0.26(55) 14.3=14.3